OpenStudy (anonymous):

Help pleaseeeee!(: The position of an object as a function of time is given by x = At2 – Bt + C, where A = 8.7 m/s2, B = 5.5 m/s, and C = 4.3 m. Find the instantaneous velocity and acceleration as functions of time. (Use the following as necessary: t.) V=_______ a=_______

OpenStudy (anonymous):

V=dx/dt V=2At-B a=dv/dt you can find a.

OpenStudy (anonymous):

how do i find dv and dt?

OpenStudy (anonymous):

dv/dt refers to derivative of v to t. So a=2A

OpenStudy (anonymous):

so a would be a=2(8.7)m/s^2 ? a=17.4 m/s^2?

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

\[v = \frac{dx}{dt}\] and \[a = \frac{d^2x}{dt^2}\] that's it.

OpenStudy (loser66):

for v, take the first derivative of x , then, plug the value of letter in. for a, take second derivative of x, then ,plug the value of letter in. Done.

OpenStudy (loser66):

got me?

OpenStudy (anonymous):

so far yes. but how do i obtain the values of dx,dt, d^2x and dt^2? @Loser66

OpenStudy (loser66):

don't you know how to take derivative?

OpenStudy (anonymous):

what am i taking the derivative of? x=At^2-Bt+C?

OpenStudy (loser66):

yes, it is.

OpenStudy (loser66):

you see, x =.... is a function of t or you can write it as x(t)

OpenStudy (anonymous):

x(t)=At^2-Bt+C ? Sorry I'm a bit confused

OpenStudy (loser66):

which grade are you? college? highschool?

OpenStudy (anonymous):

college, and I'm struggling badly /; with this and another problem

OpenStudy (loser66):

how about calculus?

OpenStudy (loser66):

I have to know where are you to give out the appropriate instruction.

OpenStudy (anonymous):

I took calculus, just trying to recall thing from it..

OpenStudy (loser66):

ok, (t^3)' =?

OpenStudy (anonymous):

3t^2?

OpenStudy (loser66):

yes, so, (At^2)' =?

OpenStudy (anonymous):

2At?

OpenStudy (loser66):

hahaha.. dan, guide him, please, don't just give out the answer

OpenStudy (dan815):

v=d/dt(f(t))

OpenStudy (anonymous):

i'm a her.

OpenStudy (dan815):

do you know v=d/t formula

OpenStudy (dan815):

speed my definition is distance travelled per unit time

OpenStudy (anonymous):

velocity= distance/time

OpenStudy (dan815):

velocity=displacement/time but okay

OpenStudy (dan815):

lets move on!

OpenStudy (dan815):

now you have a function here f(t)=x

OpenStudy (dan815):